# Leveraging Representations from Intermediate Encoder-blocks for Synthetic Image Detection

### v1

@@ Abstract
The recently developed and publicly available synthetic image generation methods and services make it possible to create extremely realistic imagery on demand, raising great risks for the integrity and safety of online information. State-of-the-art Synthetic Image Detection (SID) research has led to strong evidence on the advantages of feature extraction from foundation models. However, such extracted features mostly encapsulate high-level visual semantics instead of fine-grained details, which are more important for the SID task. On the contrary, shallow layers encode low-level visual information. In this work, we leverage the image representations extracted by intermediate Transformer blocks of CLIP’s image-encoder via a lightweight network that maps them to a learnable forgery-aware vector space capable of generalizing exceptionally well. We also employ a trainable module to incorporate the importance of each Transformer block to the final prediction. Our method is compared against the state-of-the-art by evaluating it on 20 test datasets and exhibits an average +10.6% absolute performance improvement. Notably, the best performing models require just a single epoch for training ($\sim$8 minutes). Code available at GitHub.
Keywords synthetic image detection  $\cdot$ AI-generated image detection  $\cdot$ pre-trained intermediate features
@@ 1 Introduction
The recent advancements in the field of synthetic content generation is disrupting digital media communications, and pose new societal and economic risks [1]. Generative Adversarial Networks (GAN) [2], the variety of successor generative models [3] and the latest breed of Diffusion models [4] are capable of producing highly realistic images that humans are mostly incapable of distinguishing from real ones [5], and lead to grave risks ranging from fake pornography to hoaxes, identity theft, and financial fraud [6]. Therefore, it becomes extremely challenging to develop reliable synthetic content detection methods that can keep up with the latest generative models.
@@
Recently, the utilization of features extracted by foundation models, such as CLIP [17], has been found to yield surprisingly high performance in the SID task, with minimal training requirements [16]. However, this level of performance is associated with features extracted from the last model layer, which are known to capture high-level semantics. Instead, low-level image features deriving from intermediate layers, which are known to be highly relevant for synthetic image detection [18, 15], have not been explored to date, despite their potential to further boost performance.
To address the SID task more effectively, we propose the RINE model by leveraging Representations from Intermediate Encoder-blocks of CLIP. Specifically, we collect the image representations provided by intermediate Transformer blocks that carry low-level visual information and project them with learnable linear mappings to a forgery-aware vector space. Additionally, a Trainable Importance Estimator (TIE) module is used to incorporate the impact of each intermediate Transformer block in the final prediction. We train our models only on images generated by ProGAN (our experiments consider 1, 2, and 4 object classes for training) and evaluate it on 20 test datasets including images generated from GAN, Diffusion as well as other (deepfake, low-level vision, perceptual loss, DALL-E) generative models, surpassing by an average +10.6% the state-of-the-art performance. Figure 1 provides a comparison between the feature space of RINE and that of state-of-the-art methods, for GAN, Diffussion and real images, showcasing the discriminating ability of the proposed representation. Also, it is noteworthy that (i) we reach this level of performance with only 6.3M learnable parameters, (ii) trained only for 1 epoch, which requires $\sim$8 minutes, and (iii) we surpass the state-of-the-art by +9.4%, when training only on a single ProGAN object class, while the second-to-best model uses 20.
The paper is structured as follows. Section 2 reviews the related literature. Section 3 describes the proposed method. Section 4 presents the experimental setup. Section 5 illustrates the results. Section 6 outlines the conclusions.
@@ 2 Related work
Synthetic Image Detection (SID) has recently emerged as a special but increasingly interesting field of DeepFake or synthetic media detection (term often used as an umbrella term for all kinds of synthetic or manipulated content) [6, 7]. Early studies investigated fingerprints left by GAN generators, similarly to the fingerprints left by cameras [19], letting them not only to distinguish between real and synthetic images but also to attribute synthetic images to specific GAN generators [20, 21]. Other studies focused on GAN-based image-to-image translation detection [22] and face manipulation detection [23], but still the considered evaluation sets contained images generated from the same generators as the ones used for training, neglecting to assess generalizability. Such approaches have been shown to struggle with detecting images coming from unseen generators [8].
@@
Subsequent works focused on alleviating this issue through patch-based predictions [10], auto-encoder based architectures [8] and co-occurrence matrices computed on the RGB image channels [24]. Next, Wang et al. [9] proposed a simple yet effective approach, based on appropriate pre- and post-processing as well as data augmentation, which is capable of generalizing well to many GAN generators by simply training on ProGAN images [25]. Additionally, traces of GAN generators in the frequency domain were exploited for effectively addressing the SID task [26, 11], while the use of frequency-level perturbation maps was proposed in [27] to force the detector to ignore domain-specific frequency-level artifacts. Image gradients estimated by a pre-trained CNN model have been shown to produce artifacts that can generalize well to unseen GAN generated images [28]. Removing the down-sampling operation from the first ResNet’s layer, that potentially eliminates synthetic traces, together with intense augmentation has shown promising results [15]. Another approach to improve generalization has been to select training images based on their perceptual quality; this was shown to be effective especially in cross-domain generalization settings [29].
@@
Pre-trained features extracted by a foundation model were demonstrated to be surprisingly effective for the SID task, generalizing equally well on GAN and Diffusion model generated data [16]. Motivated by this success, we hypothesize that further performance gains are possible by leveraging representations from intermediate layers, which carry low-level visual information, in addition to representations from the final layer that primarily carry high-level semantic information. Our approach is different from deep layer aggregation that was explored in several computer vision tasks [30], as well as for deepfake detection [31], as we process the frozen layers’ outputs of a large-scale foundation model, while deep layer aggregation is a training technique that concatenates layer outputs to produce predictions during the training of a CNN model.
@@
Finally, the performance drop of synthetic image detectors in real-world conditions typically performed when uploading content to social media (e.g., cropping and compression) has been analysed in [22, 15]. Our method is also effective in such conditions as demonstrated by the experimental results in Section 5.6.
@@ 3 Methodology
Figure 2 illustrates the main elements of the RINE architecture. We build upon the intermediate layer outputs of CLIP’s image encoder in order to train a lightweight SID model.
Figure 2: The RINE architecture. A batch of $b$ images is processed by CLIP’s image encoder. The concatenation of the $n$ $d$-dimensional CLS tokens (one from each intermediate stage, i.e., Transformer block) is first projected and then multiplied with the trainable importance scores of the intermediate stages, estimated by the Trainable Importance Estimator (TIE) module. Summation across the second dimension results in one feature vector per image. Finally, after the second projection and the consequent classification head modules, two loss functions are computed. The binary cross-entropy $\mathfrak{L}_{CE}$ directly optimizes the real vs. fake objective, while the contrastive loss $\mathfrak{L}_{Cont.}$ assists the training by forming a dense feature vector cluster per class.
@@ 3.1 Representations from Intermediate Encoder-blocks
Let us consider a batch of $b$ input images $\mathbf{X}\in\mathbb{R}^{b\times 3\times w\times h}$, with $w$ width and $h$ height. The images are first reshaped into a sequence of $p$ flattened image patches $\mathbf{X}_{p}\in\mathbb{R}^{b\times p\times(P^{2}\cdot 3)}$, where $P$ denotes patch side length and $p=w\cdot h/P^{2}$. Then, a $d$-dimensional linear mapping is considered on top of $\mathbf{X}_{p}$, the learnable $d$-dimensional CLS token is concatenated to the projected sequence, and positional embeddings are added to all $p+1$ tokens in order to construct the image encoder’s input, $\mathbf{Z}_{0}\in\mathbb{R}^{b\times(p+1)\times d}$. CLIP’s image encoder [17], being a standard Vision Transformer (ViT) [32], consists of $n$ successive Transformer encoder blocks [33], each processing the output of the previous block as shown in Equation 1:
@@
where $l$=1,…,n denotes the Transformer block’s index, $\mathbf{Z}_{l}\in\mathbb{R}^{b\times(p+1)\times d}$ denotes the $l$th Transformer block’s output, MSA denotes Multi-head Softmax Attention [33], LN denotes Layer Normalization [34], and MLP denotes a Multi-Layer Perceptron with two GELU [35] activated linear layers of $4\cdot d$ and $d$ number of units, respectively.
Finally, we define the tensor $\mathbf{K}$ containing the Representations from Intermediate Encoder-blocks (RINE), as the concatenation of the $n$ CLS tokens stemming from each of the $n$ Transformer blocks:
@@
where $\oplus$ denotes concatenation, and $\mathbf{Z}_{l}^{\text{[0]}}\in\mathbb{R}^{b\times 1\times d}$ denotes the CLS token from the output of Transformer block $l$. We keep CLIP, which produces these features, frozen during training and we use $\mathbf{K}$ to construct discriminative features for the SID task.
@@ 3.2 Trainable modules
The extracted representations $\mathbf{K}$ are processed by a projection network ($\mathfrak{Q}_{1}$ in Figure 2):
@@
Additionally, each of the $d^{\prime}$ learned features can be more relevant for the SID task at various processing stages, thus we employ a Trainable Importance Estimator (TIE) module to adjust their impact to the final decision. More precisely, we consider a randomly-initialized learnable variable $\mathbf{A}=\{\alpha_{lk}\}\in\mathbb{R}^{n\times d^{\prime}}$ the elements of which will estimate the importance of feature $k$ at processing stage (i.e., Transformer block) $l$. This is then used to construct one feature vector per image, as a weighted average of features across processing stages:
$\tilde{\mathbf{K}}=\sum_{l}^{n}\mathcal{S}(\mathbf{A})\otimes\mathbf{K}_{q}$ (4)
where $\mathcal{S}$ denotes the Softmax activation function acting across the first dimension of $\mathbf{A}$, $\otimes$ denotes broadcasted Hadamard product, and summation is conducted across the dimension of Transformer blocks.
Finally, a second projection network ($\mathfrak{Q}_{2}$ in Figure 2) with the same architecture as the first one (cf. Equation 3) takes as input $\tilde{\mathbf{K}}$ and outputs $\tilde{\mathbf{K}}_{q}$, which is consequently processed by the classification head that predicts the final output (real/fake). The classification head consists of two $d^{\prime}\times d^{\prime}$ ReLU-activated dense layers and one $d^{\prime}\times 1$ dense layer which produces the logits.
@@ 3.3 Objective function
We consider the combination of two objective functions to optimize the parameters of the proposed model. The first is the binary cross-entropy loss [38], here denoted as $\mathfrak{L}_{CE}$, which measures the classification error and directly optimizes the main SID objective. The second is a contrastive loss function, specifically the Supervised Contrastive Learning loss (SupContrast) [39], which we here denote as $\mathfrak{L}_{Cont.}$, and is considered in order to assist the training process by bringing closer the feature vectors inside $\tilde{\mathbf{K}}_{q}$ that share targets and move apart the rest. We combine the two objective functions by a tunable factor $\xi$, as shown in Equation 5:
@@ 4.2 Implementation details
The training of RINE is conducted with batch size 128 and learning rate 1e-3 for only 1 epoch using the Adam optimizer. To assess the existence of potential benefit from further training we also conduct a set of experiments with 3, 5, 10, and 15 epochs (cf. Section 5.4). For each of the 3 training settings (1-class, 2-class, 4-class), we consider a hyperparameter grid to obtain the best performance, namely $\xi\in\{0.1,0.2,0.4,0.8\}$, $q\in\{1,2,4\}$, and $d^{\prime}\in~{}\{128,256,512,1024\}$. Additionally, we consider two CLIP variants, namely ViT-B/32 and L/14, for the extraction of representations. The models’ hyperparameters that result in the best performance (presented in Section 5) are illustrated in Table 1, along with the corresponding number of trainable parameters, and required training duration. For the training images, following the best practice in related work, we apply Gaussian blurring and $\mathtt{JPEG}$ compression with probability 0.5, then random cropping to 224$\times$224, and finally random horizontal flip with probability 0.5. The validation and test images are only center-cropped at 224$\times$224. Resizing is omitted both during training and testing as it is known to eliminate synthetic traces [15]. All experiments are conducted using one NVIDIA GeForce RTX 3090 Ti GPU.
@@ 4.3 State-of-the-art methods
    DMID [15]: A ResNet-50 [54] without the down-sampling in the first layer and intense augmentation trained on the SID task.
@@
For all methods except FrePGAN, we compute performance metrics on the evaluation datasets presented in Section 4.1, using the publicly available checkpoints provided by their official repositories. For FrePGAN, that does not provide publicly available code and models, we present the scores that are reported in their paper [27].

### v2

# Christos Koutlis, Symeon Papadopoulos CERTH-ITI Thessaloniki, Greece

@@ Abstract
The recently developed and publicly available synthetic image generation methods and services make it possible to create extremely realistic imagery on demand, raising great risks for the integrity and safety of online information. State-of-the-art Synthetic Image Detection (SID) research has led to strong evidence on the advantages of feature extraction from foundation models. However, such extracted features mostly encapsulate high-level visual semantics instead of fine-grained details, which are more important for the SID task. On the contrary, shallow layers encode low-level visual information. In this work, we leverage the image representations extracted by intermediate Transformer blocks of CLIP’s image-encoder via a lightweight network that maps them to a learnable forgery-aware vector space capable of generalizing exceptionally well. We also employ a trainable module to incorporate the importance of each Transformer block to the final prediction. Our method is compared against the state-of-the-art by evaluating it on 20 test datasets and exhibits an average +10.6% absolute performance improvement. Notably, the best performing models require just a single epoch for training ($\sim$8 minutes). Code available at https://github.com/mever-team/rine.
Keywords synthetic image detection  $\cdot$ AI-generated image detection  $\cdot$ intermediate representations
@@ 1 Introduction
The advancements in the field of synthetic content generation disrupt digital media communications, and pose new societal and economic risks [1]. Generative Adversarial Networks (GAN) [2], the variety of successor generative models [3] and the latest breed of Diffusion models [4] are capable of producing highly realistic images that often deceive humans [5], and lead to grave risks ranging from fake pornography to hoaxes, identity theft, and financial fraud [6]. Therefore, it becomes extremely challenging to develop reliable synthetic content detection methods that can keep up with the latest generative models.
@@
Recently, the utilization of features extracted by foundation models, such as CLIP [17], has been found to yield surprisingly high performance in the SID task, with minimal training requirements [16]. However, these features are extracted from the last model layer, known to capture high-level semantics. Instead, low-level image features deriving from intermediate layers, which are known to be highly relevant for synthetic image detection [18, 15], have not been explored to date, despite their potential to further boost performance.
To address the SID task more effectively, we propose the RINE model by leveraging Representations from Intermediate Encoder-blocks of CLIP. Specifically, we collect the image representations provided by intermediate Transformer blocks that carry low-level visual information and project them with learnable linear mappings to a forgery-aware vector space. Additionally, a Trainable Importance Estimator (TIE) module is used to incorporate the impact of each intermediate Transformer block in the final prediction. We train our models only on images generated by ProGAN (our experiments consider 1, 2, and 4 object classes for training) and evaluate it on 20 test datasets including images generated from GAN, Diffusion as well as other (deepfake, low-level vision, perceptual loss, DALL-E) generative models, surpassing by an average +10.6% the state-of-the-art performance. Figure˜1 provides a comparison between the feature space of RINE and that of state-of-the-art methods, for GAN, Diffussion and real images, showcasing the discriminating ability of the proposed representation. Also, it is noteworthy that (i) we reach this level of performance with only 6.3M learnable parameters, (ii) trained only for 1 epoch, which requires $\sim$8 minutes, and (iii) we surpass the state-of-the-art by +9.4%, when training only on a single ProGAN object class, while the second-to-best model uses 20.
@@ 2 Related Work
Synthetic Image Detection (SID) has recently emerged as a special but increasingly interesting field of DeepFake [^1] Often used as an umbrella term for all kinds of synthetic or manipulated content or synthetic media detection [6, 7]. Early studies investigated fingerprints left by GAN generators, similarly to the fingerprints left by cameras [19], letting them not only to distinguish between real and synthetic images but also to attribute synthetic images to specific GAN generators [20, 21]. Other studies focused on GAN-based image-to-image translation detection [22] and face manipulation detection [23], but still the considered evaluation sets contained images generated from the same generators as the ones used for training, neglecting to assess generalizability. Such approaches have been shown to struggle with detecting images coming from unseen generators [8].
@@
Subsequent works focused on alleviating this issue through patch-based predictions [10], auto-encoder based architectures [8] and co-occurrence matrices computed on the RGB image channels [24]. Next, Wang et al. [9] proposed a simple yet effective approach, based on appropriate pre-processing and intense data augmentation, which is capable of generalizing well to many GAN generators by simply training on ProGAN images [25]. Additionally, traces of GAN generators in the frequency domain were exploited for effectively addressing the SID task [26, 11], while the use of frequency-level perturbation maps was proposed in [27] to force the detector to ignore domain-specific frequency-level artifacts. Image gradients estimated by a pre-trained CNN model have been shown to produce artifacts that can generalize well to unseen GAN generated images [28]. Removing the down-sampling operation from the first ResNet’s layer, that potentially eliminates synthetic traces, together with intense augmentation has shown promising results [15]. Another approach to improve generalization has been to select training images based on their perceptual quality; this was shown to be effective especially in cross-domain generalization settings [29].
@@
Representations extracted by foundation models are surprisingly effective on SID, generalizing equally well on GAN and Diffusion model generated data [16]. Motivated by this success, we hypothesize that further performance gains are possible by leveraging representations from intermediate layers, which carry low-level visual information, in addition to representations from the final layer that primarily carry high-level semantic information. Our approach is different from deep layer aggregation that was explored in several computer vision tasks [30], as well as for deepfake detection [31], as we process the frozen layers’ outputs of a foundation model, while deep layer aggregation is a training technique that concatenates layer outputs to produce predictions during the training of a CNN.
@@
Finally, the performance drop of synthetic image detectors in real-world conditions typically performed when uploading content to social media (e.g., cropping and compression) has been analysed in [22, 15]. Our method is also effective in such conditions as demonstrated by the experimental results in section˜5.6.
@@ 3 Methodology
Figure 2: The RINE architecture. A batch of $b$ images is processed by CLIP’s image encoder. The concatenation of the $n$ $d$-dimensional CLS tokens (one from each Transformer block) is first projected and then multiplied with the blocks’ scores, estimated by the Trainable Importance Estimator (TIE) module. Summation across the second dimension results in one feature vector per image. Finally, after the second projection and the consequent classification head modules, two loss functions are computed. Binary cross-entropy $\mathfrak{L}_{CE}$ directly optimizes SID, while the contrastive loss $\mathfrak{L}_{Cont.}$ assists the training by forming a dense feature vector cluster per class.
@@ 3.1 Representations from Intermediate Encoder-Blocks
Figure˜2 illustrates the RINE architecture. Let us consider a batch of $b$ input images $\mathbf{X}\in\mathbb{R}^{b\times 3\times w\times h}$, with $w$ width, and $h$ height, and targets $\mathbf{Y}=\{y_{i}\}_{i=1}^{b}\in\mathbb{R}^{b}$, with $y_{i}\in\{0,1\}$. The images are reshaped into a sequence of $p$ flattened patches $\mathbf{X}_{p}\in\mathbb{R}^{b\times p\times(P^{2}\cdot 3)}$, where $P$ denotes patch side length and $p=w\cdot h/P^{2}$. Then, $\mathbf{X}_{p}$ is linearly projected to $d$ dimensions, the learnable $d$-dimensional CLS token is concatenated to the projected sequence, and positional embeddings are added to all $p+1$ tokens in order to construct the image encoder’s input, $\mathbf{Z}_{0}\in\mathbb{R}^{b\times(p+1)\times d}$. CLIP’s image encoder [17], being a Vision Transformer (ViT) [32], consists of $n$ successive Transformer encoder blocks [33], each processing the output of the previous block as shown in eq.˜1:
@@
where $l$=1,…,n denotes the block’s index, $\mathbf{Z}_{l}\in\mathbb{R}^{b\times(p+1)\times d}$ denotes the $l$th block’s output, MSA denotes Multi-head Softmax Attention [33], LN denotes Layer Normalization [34], and MLP denotes a Multi-Layer Perceptron with two GELU [35] activated linear layers of $4\cdot d$ and $d$ number of units, respectively.
Then, we define the Representations from Intermediate Encoder-blocks (RINE) $\mathbf{K}$, as the concatenation of CLS tokens from the corresponding $n$ blocks:
@@
where $\oplus$ denotes concatenation, and $\mathbf{Z}_{l}^{\text{[0]}}\in\mathbb{R}^{b\times 1\times d}$ denotes the CLS token from the output of block $l$. We keep CLIP, which produces these features, frozen during training and use $\mathbf{K}$ to construct discriminative features for the SID task.
@@ 3.2 Trainable Modules
The representations $\mathbf{K}$ are processed by a projection network ($\mathfrak{Q}_{1}$ in fig.˜2):
@@
Additionally, each of the $d^{\prime}$ learned features can be more relevant for the SID task at various processing stages, thus we employ a Trainable Importance Estimator (TIE) module to adjust their impact to the final decision. More precisely, we consider a randomly-initialized learnable variable $\mathbf{A}=\{\alpha_{lk}\}\in\mathbb{R}^{n\times d^{\prime}}$ the elements of which will estimate the importance of feature $k$ at processing stage (i.e., Transformer block) $l$. This is used to construct one feature vector per image $i$, through weighted averages of features $k$ across processing stages $l$:
$\tilde{\mathbf{K}}^{(ik)}=\sum_{l}^{n}\mathcal{S}(\mathbf{A})^{(lk)}\cdot\mathbf{K}_{q}^{(ilk)}$ (4)
where (⋅) denotes tensor indices, $\mathbf{K}_{q}=\{\mathbf{K}_{q}^{(ilk)}\}\in\mathbb{R}^{b\times n\times d^{\prime}}$ is the output of projection $\mathfrak{Q}_{1}$, and $\mathcal{S}$ denotes the Softmax activation function acting across the first dimension of $\mathbf{A}$.
Finally, a second projection network ($\mathfrak{Q}_{2}$ in fig.˜2) with the same architecture as the first one (cf. eq.˜3) takes as input $\tilde{\mathbf{K}}\in\mathbb{R}^{b\times d^{\prime}}$ and outputs $\tilde{\mathbf{K}}_{q}\in\mathbb{R}^{b\times d^{\prime}}$, which is consequently processed by the classification head that predicts the final output $\hat{y}_{i}$ (probability to be fake). The classification head consists of two $d^{\prime}\times d^{\prime}$ ReLU-activated dense layers and one $d^{\prime}\times 1$ dense layer which produces the logits.
@@ 3.3 Objective Function
We consider the combination of two objective functions to optimize the parameters of the proposed model. The first is the binary cross-entropy loss [38], which measures the classification error and directly optimizes the SID objective:
$\mathfrak{L}_{CE}=-\sum_{i=1}^{b}y_{i}\text{log}\hat{y}_{i}+(1-y_{i})\cdot\text{log}(1-\hat{y}_{i})$
The second is the Supervised Contrastive Learning loss (SupContrast) [39], which is considered in order to assist the training process by bringing closer the feature vectors inside $\tilde{\mathbf{K}}_{q}$ that share targets and move apart the rest, defined as:
$\mathfrak{L}_{Cont.}=-\sum_{i=1}^{b}\frac{1}{G(i)}\sum_{g\in G(i)}\text{log}\frac{\text{exp}(\mathbf{z}_{i}\cdot\mathbf{z}_{g}/\tau)}{\sum_{a\in A(i)}\text{exp}(\mathbf{z}_{i}\cdot\mathbf{z}_{a}/\tau)}$
where $A(i)=\{1,\dots,i-1,i+1,\dots,b\}$, $G(i)=\{g\in A(i):y_{g}=y_{i}\}$, $\mathbf{z}_{j}=\{\tilde{\mathbf{K}}_{q}^{(jk)}\}_{k=1}^{d^{\prime}}\in\mathbb{R}^{d^{\prime}}$, the $\cdot$ denotes dot product, and $\tau$ is a temperature parameter. We combine the two loss functions by a tunable factor $\xi$, as shown in eq.˜5:
@@ 4.2 Implementation Details
The training of RINE is conducted with batch size 128 and learning rate 1e-3 for only 1 epoch using the Adam optimizer. To assess the existence of potential benefit from further training we also conduct a set of experiments with 3, 5, 10, and 15 epochs (cf. section˜5.4). For each of the 3 training settings (1-class, 2-class, 4-class), we consider a hyperparameter grid to obtain the best performance, namely $\xi\in\{0.1,0.2,0.4,0.8\}$, $q\in\{1,2,4\}$, and $d^{\prime}\in~\{128,256,512,1024\}$. Additionally, we consider two CLIP variants, namely ViT-B/32 and L/14, for the extraction of representations. The models’ hyperparameters that result in the best performance (presented in section˜5) are illustrated in table˜1, along with the corresponding number of trainable parameters, and required training duration. For the training images, following the best practice in related work, we apply Gaussian blurring and $\mathtt{JPEG}$ compression with probability 0.5, then random cropping to 224$\times$224, and finally random horizontal flip with probability 0.5. The validation and test images are only center-cropped at 224$\times$224. Resizing is omitted both during training and testing as it is known to eliminate synthetic traces [15]. All experiments are conducted using one NVIDIA GeForce RTX 3090 Ti GPU.
@@ 4.3 State-of-the-Art Methods
    DMID [15]: A ResNet-50 [54] without down-sampling in the first layer is trained on the SID task using intense augmentation.
@@
For all methods except FrePGAN, we compute performance metrics on the evaluation datasets presented in section˜4.1, using the publicly available checkpoints provided by their official repositories. For FrePGAN, that does not provide publicly available code and models, we present the scores that are reported in their paper [27].